# Urban Anomalies: A Simulated Human Mobility Dataset with Injected Anomalies

### v1

@@ 1. Introduction
Anomaly detection algorithms have been extensively studied in the literature (Lee et al., 2008; Yu et al., 2017; Zhang et al., 2023; Lv et al., 2017). However, the absence of a standardized benchmark dataset remains a critical challenge in this domain (Campos et al., 2016). Researchers often rely on various real-world datasets to evaluate their algorithms, but these datasets typically lack ground truth information. The UCI GPS database comprises 603 car driving trajectories with over 5,317 points, where each trajectory contains more than 2,000 points, making it relatively sparse (Cruz et al., 2015). Similarly, the Geolife trajectory database, collected by Microsoft Research Asia, includes 17,621 trajectories with over 152,241 points, with each trajectory exceeding 5,000 points, also contributing to sparsity (Zheng et al., 2011). In contrast, the T-Drive trajectory database, also from Microsoft Research Asia, represents a dense dataset generated by over 33,000 taxis in a period of 3 months which does not have ground truth for anomaly detection tasks (Yuan et al., 2010). Furthermore, the field would greatly benefit from the development of large-scale, simulated datasets that incorporate human mobility, check-ins, and social networks, such as those generated using location-based social network simulation models (Amiri et al., 2023).
In this paper, we leverage an existing Pattern of Life Simulation model (Züfle et al., 2023; Kim et al., 2020) to simulate and identify anomaly behaviors in agent-based environments. This simulation roots the behavior of agents on Maslowian Needs (Maslow, 1943), such as Physiological Needs (such as Food Need), Safety Needs (such as Financial Need), and Love Needs (such as a need to maintain friendships). Details on these needs and their prioritization that leads to agent actions and behavior are detailed in (Züfle et al., 2023). To inject anomalous behavior into this simulation, we define anomalies inspired by the definition of Outliers provided by Hawkins (Hawkins, 1980), which describes an outlier as “an observation that deviates significantly from other observations, raising suspicions of being generated by a different mechanism” (Barnett et al., 1994). To model anomalous agent behavior, we inject four types of anomalous behavior:
@@
For each of the aforementioned types of anomalous behaviors, and for each type of injection mechanism, we provide data resulting from a simulation of 1000 agents, each having four weeks of normal behavior (without any anomalies) and having at least four weeks of anomalous behavior (during which anomalies are injected as described above). We hope that these datasets will allow scholar to investigate novel approaches to identify anomalous behavior in simulated human mobility data.
@@
Mobility Information Urban Anomalies Dataset
@@ 2. Related Work
A critical aspect of working with GPS data is the detection of anomalies, especially location-based anomalies. This process is fundamental to various algorithms employed in both industrial applications and research studies. Accurate identification and management of such anomalies help ensure the reliability and precision of models that rely on spatial data, which can greatly affect the outcomes of analyses ranging from navigation systems to geographical research (Kriegel et al., 2010). Anomaly detection algorithms have been the focus of extensive research in the literature (Lee et al., 2008; Yu et al., 2017; Zhang et al., 2023, 2024; Lv et al., 2017). Despite this, a significant challenge persists in the form of the absence of a standardized benchmark dataset for the field (Campos et al., 2016). The lack of such a benchmark complicates the comparison and evaluation of different algorithms, as researchers are often compelled to rely on a variety of real-world datasets that typically lack ground truth information. This absence of consistent, labeled data makes it difficult to objectively assess the performance and generalizability of anomaly detection methods across different contexts and applications.
@@
A comprehensive survey on trajectory analysis within the video surveillance domain, with a particular emphasis on trajectory anomaly detection, is presented in (Belhadi et al., 2020). This survey examines the extensive research conducted in various industrial applications, such as maritime, urban transportation, and networking. It offers an in-depth overview of existing trajectory anomaly detection algorithms, categorizing them according to different criteria, including application, output, and algorithmic levels. The paper also identifies the key challenges and open issues within the field. Furthermore, it provides an extensive list of publicly available datasets used for trajectory analysis, including notable examples like the UCI GPS database, the Geolife trajectory database, and the Manhattan taxi dataset.
@@ 3.2. Simulation of Anomalies
To simulate the data, we employ a Patterns of Life Simulation and inject a set of agents specifically designed to exhibit anomaly behavior. These agents the source code of the simulation is available at https://github.com/onspatial/sigspatial2024-anomaly-dataset
@@ 4.2. Datasets Analysis
This section provides the basic analysis of the generated datasets, in order to show the existence and capability to distinguish the generated anomaly, as well as the specific patterns in the datasets with the infectious disease model involved.

### v2

@@ 1. Introduction
Anomaly detection algorithms have been extensively studied in the literature (Lee et al., 2008; Yu et al., 2017; Zhang et al., 2023; Lv et al., 2017). However, the absence of a standardized benchmark dataset remains a critical challenge in this domain (Campos et al., 2016; Stanford et al., 2024). Researchers often rely on various real-world datasets to evaluate their algorithms, but these datasets typically lack ground truth information. The UCI GPS database comprises 603 car driving trajectories with over 5,317 points, where each trajectory contains more than 2,000 points, making it relatively sparse (Cruz et al., 2015). Similarly, the Geolife trajectory database, collected by Microsoft Research Asia, includes 17,621 trajectories with over 152,241 points, with each trajectory exceeding 5,000 points, also contributing to sparsity (Zheng et al., 2011). In contrast, the T-Drive trajectory database, also from Microsoft Research Asia, represents a dense dataset generated by over 33,000 taxis in a period of 3 months which does not have ground truth for anomaly detection tasks (Yuan et al., 2010). Furthermore, the field would greatly benefit from the development of large-scale, simulated datasets that incorporate human mobility, check-ins, and social networks, such as those generated using location-based social network simulation models (Amiri et al., 2023).
In this paper, we leverage an existing Pattern of Life Simulation model (Züfle et al., 2023; Kim et al., 2020) to simulate and identify anomaly behaviors in agent-based environments. This simulation roots the behavior of agents on Maslowian Needs (Maslow, 1943), such as Physiological Needs (e.g. Food Need), Safety Needs (e.g. Financial Need), and Love Needs (e.g. a need to maintain friendships). Details on these needs and their prioritization that leads to agent actions and behavior are detailed in (Züfle et al., 2023). To add anomalous behavior into this simulation, we define anomalies as deviations from normal patterns that suggest they were generated by a different underlying mechanism (Hawkins, 1980; Barnett et al., 1994). We inject four types of anomalies at three different intensity levels to simulate abnormal patterns of life, using three distinct mechanisms to inject these anomalies. To model anomalous agent behavior, we inject four types of anomalous behavior:
@@
For each of the aforementioned types of anomalous behaviors, and for each type of injection mechanism, we provide data resulting from a simulation of 1000 agents, each having four weeks of normal behavior (without any anomalies) and having at least four weeks of anomalous behavior (during which anomalies are injected as described above). We hope that these datasets will allow scholars to investigate novel approaches to identify anomalous behavior in simulated human mobility data.
@@
Mobility Information Urban Anomalies
@@ 2. Related Work
A critical aspect of working with GPS data is the detection of anomalies, especially location-based anomalies. This process is fundamental to various algorithms employed in both industrial applications and research studies. Accurate identification and management of such anomalies help ensure the reliability and precision of models that rely on spatial data, which can greatly affect the outcomes of analyses ranging from navigation systems to geographical research (Kriegel et al., 2010). Anomaly detection algorithms have been the focus of extensive research in the literature (Lee et al., 2008; Yu et al., 2017; Zhang et al., 2023, 2024; Lv et al., 2017; Liu et al., 2024). Despite this, a significant challenge persists in the form of the absence of a standardized benchmark dataset for the field (Campos et al., 2016). The lack of such a benchmark complicates the comparison and evaluation of different algorithms, as researchers are often compelled to rely on a variety of real-world datasets that typically lack ground truth information. This absence of consistent, labeled data makes it difficult to objectively assess the performance and generalizability of anomaly detection methods across different contexts and applications.
@@
A comprehensive survey on trajectory analysis within the video surveillance domain, with a particular emphasis on trajectory anomaly detection, is presented in (Belhadi et al., 2020). This survey examines the extensive research conducted in various industrial applications, such as maritime, urban transportation, and networking. It offers an in-depth overview of existing trajectory anomaly detection algorithms, categorizing them according to different criteria, including application, output, and algorithmic levels. The paper also identifies the key challenges and open issues within the field. Furthermore, it provides an extensive list of publicly available datasets used for trajectory analysis, including notable examples like the UCI GPS, the Geolife trajectory, and the Manhattan taxi datasets.
@@ 3.2. Simulation of Anomalies
To simulate the data, we employ a Patterns of Life Simulation and inject a set of agents specifically designed to exhibit anomaly behavior. The source code of the simulation is available at https://github.com/onspatial/sigspatial2024-anomaly-dataset
@@ 4.2. Datasets Analysis
This section provides the basic analysis of the generated datasets for three types of anomalies: hunger, work, and social anomalies to show the existence and capability to distinguish the generated anomaly, as well as the specific patterns in the datasets with the infectious disease model involved. Notably, the interest anomalies need an in-depth social network analysis to be distinguished which is not included in the basic analysis.